OpenStudy (mackenzie2013):

Write the fractional equivalent (in reduced form) of each number. 0.1 0.16(line over six) 0.6(line over six) 0.6

OpenStudy (mackenzie2013):

Please explain .. I dont get this

OpenStudy (anonymous):

a.1/10 b.16/100 on reduccing, 8/50=4\25 c.6/10, on reducing, 3/5 d.

OpenStudy (anonymous):

0.1=1/10 .16=4/25 .6=2/3 .6=3/5

OpenStudy (mackenzie2013):

Thank you :)

OpenStudy (anonymous):

YW